OpenStudy (anonymous):

Write the equation of the line that passes through (2, 4) and (1, –3) in slope-intercept form. y = 7x – 10 y = 7x + 10 y = 7x + 18 y = 7x – 18

OpenStudy (anonymous):

plug in one point to y=7x+b. so 4=7(2)+b

OpenStudy (anonymous):

y=7x+ your answer from plugging in

OpenStudy (anonymous):

The equation of the gradient of this line is \[m=\frac{ y _{1}-y_{2} }{ x_{1}-x_{2} }\] So your two points (2,4) (1,-3) can be used to find the gradient, we see that it is \[m=\frac{ 4--3 }{ 2-1 }\] which gives m=7. Now the equation of a straight line is given by \[y-y_{1}=m(x-x_{1})\] which you can use either of your two coordinates, they will both give the same answer. Using (2,4) as it is easier gives us. \[y-4=7(x-2)\]\[y-4=7x-14\]\[y=7x-10\] Which is your final answer :)